40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jun 19 17:06:50 2020, gautam, Update, LSC, What should the short-term commissioning goals be? CARM_superBoost.pdf
    Reply  Wed Jun 24 17:20:16 2020, gautam, Update, LSC, What should the short-term commissioning goals be? 
       Reply  Tue Jun 30 10:59:16 2020, gautam, Update, LSC, Three sensing matrices PRMI_1f_20200625sensMat.pdfPRFPMI_20200629sensMat.pdfPRFPMI_20200629sensMat_wOffset.pdf
          Reply  Mon Jul 6 17:41:19 2020, gautam, Update, LSC, An LSC puzzler LSCsignals.jpg
Message ID: 15419     Entry time: Fri Jun 19 17:06:50 2020     Reply to this: 15427
Author: gautam 
Type: Update 
Category: LSC 
Subject: What should the short-term commissioning goals be? 

Summary:

I want some input about what the short-term (next two weeks) commissioning goals should be.

Details:

Before the vacuum fracas, the locking was pretty robust. With some human servoing of the input beam, I could maintain locks for ~1 hour. My primary goals were:

  1. Transition the vertex length DoF control from 3f signals to 1f signals.
  2. Turn on some MICH-->DARM feedforward cancellation, because the noise between ~100 Hz and ~1kHz is dominated by this cross-coupling.

I didn't succeed in either so far.

  1. I find that there is poor separation of the length DoFs in the 1f sensors, which makes this transition hopeless.
    • Why should this be? I can't get any sensing matrix in Finesse to line up with what I measure in-lock.
    • One hypothesis I came up with (but haven't yet tested) is that the offsets from the 3f photodiodes are changing from time to time, which somehow changes the projection of the various DoFs onto the photodiode quadratures. 
    • The attached GIF shows the variation in the measured sensing matrix on two days - while the sensing of MICH/PRCL in the 3f photodiodes have hardly changed, they are significantly different in the 1f photodiodes. Note that the I and Q have changed for REFL11 and REFL55 between the two days because I changed the demod phase.
    • I also thought that maybe the CARM suppression isn't sufficient for REFL11 to be used as a PRCL sensor - but even after engaging a CM board SuperBoost, I was unable to realize the PRCL 3f-->1f transition, even though the CARM-->REFL11 coupling did get smaller in the measured sensing matrix (red line in the GIF). I don't think we can juice up the CARM gain much more without modifying the CM board boosts, see Attachment #1.
  2. I was able to measure the MICH CTRL --> DARM ERR transfer function with somewhat high coherence (~0.98).
    • I then used the infrastructure available in the LSC model to try and implement some cancellation, but didn't really see any effect.
    • Perhaps the TF needs to be measured with higher coherence.
    • It may also be that if I am able to successfully execute the 3f-->1f transition, the coupling gets smaller because the 1f sensing noise is lower?

I guess apart from this, we want to run the ALS scan to try and infer something about the absorption-induced thermal lens. I guess at this point, the costs outweigh the benefits in trying to bring in the SRC as well, since we will be changing the SRC config?

Attachment 1: CARM_superBoost.pdf  104 kB  Uploaded Fri Jun 19 18:40:44 2020  | Hide | Hide all
CARM_superBoost.pdf
ELOG V3.1.3-